OpenStudy (anonymous):

Simplify the expression: 1) 1/tan^2x+1 2) sec^2x-1/secx-1 3)csc^2x(1-cos^2x) 4) x+1=0 Any help is much appreciated.

OpenStudy (loser66):

\[\frac{1}{tan^2x+1}= \frac{1}{sec^2x}=cos^2x\] wonder why don't you try? trigs is fun and very easy, just playing with formula.

OpenStudy (loser66):

Iamnotalegend, Iamghostandevenaghostcandoit.

OpenStudy (anonymous):

i like it when i understand it. I just get stuck. like how did you get from sec^2x=cos^2x

OpenStudy (anonymous):

lol that was really clever

OpenStudy (loser66):